OpenStudy (anonymous):

Solve three-fourths b minus one-sixth equals one-half for b. b equals eight-ninths b equals one-half b equals one-fourth b equals four-ninths

OpenStudy (anonymous):

use equation?

OpenStudy (anonymous):

because i didnt understand the question,rewrite it

OpenStudy (anonymous):

@aajugdar this is how It comes

OpenStudy (lncognlto):

3/4(b) - 1/6 = 1/2 3/4(b) = 1/2 + 1/6 = 3/6 + 1/6 = 4/6 = 2/3 3/4(b) = 2/3 b = 2/3(4/3) = 8/9

OpenStudy (lncognlto):

You're welcome :)